12. What is the return on a portfolio containing the stocks below? (10) Stock Shares Po P 10 12 300 20 25 500 Y
Answer #1

Value of portfolio at time 0 = Nx*Px0 + Ny*Py0 = 500*10 + 300*20 = 11000

Value of portfolio at time 1 = Nx*Px1 + Ny*Py1 = 500*12 + 300*25 = 13500

Return on portfolio = (Value at time 1 - value at time 0)/value at time 0 = (13500-11000)/11000 = 22.73%